# F-91191 Gif-sur-Yvette cedex, France

### v1

@@ 5 Conclusions
We contrast here two complementary observations. On the one hand, the transformation of the NP scalars under general coordinates transformations, large or small, suggests that their leading order in the $1/|\vec{x}|$ expansion cannot be changed by diffeomorphisms that preserves the asymptotic-Minkowski nature of the spacetime. In particular, the leading-order peeling violation cannot be removed by such transformations. On the other, a BMS transformation with parameter twice that of Veneziano and Vilkovisky Veneziano:2022zwh flips the sign of the disconnected-cut contribution to the waveform Bini:2024rsy ; Elkhidir:2024izo and, in an off-shell regularization Elkhidir:2024izo , appears to completely cancel the tail-related logarithmic terms in Eq. (2.2), $iG(p_{1}\cdot k\ln(u_{1}\cdot k)^{2}/k^{2})+p_{2}\cdot k\ln(u_{2}\cdot k)^{2}/k^{2}))$. In this BMS frame, which is distinct from both the canonical and the intrinsic ones, the novel peeling violation that we identified in Sec 4.4 appears to cancel out. We leave a resolutin of this tension for future work. Possible approaches may rely e.g. on the transformation of the parameters defining the asymptotic states under BMS transformations or on the development of a better understanding of allowed coordinate choices that preserve asymptotic flatness and of symmetries in spacetimes sourced by scattering processes.
It would be fascinating if the subleading components of the metric in $|\vec{x}|^{-1}$ were experimentally measurable. Naively, these effects are enormously suppressed by the vast distance between astrophysical events and gravitational-wave observatories, making them largely formal considerations. Nevertheless, the computational techniques developed in this paper could prove useful in contexts where finite-distance effects become relevant, such as at colliders Collins:2019ozc .
The framework developed in this paper can be used to compute the scattering metric throughout spacetime (but sufficiently far from the scattering bodies for consistency with the PM expansion). This calculation amounts to evaluating exactly the Fourier transform over the external-graviton momentum of the final-state graviton one-point function. Assuming that such an exact metric were available, it should be possible to direct test of asymptotic simplicity by computing the conformal structure and demonstrating that they are not regular at null infinity.
@@ Acknowledgements.
We thank Martin Beneke, Donato Bini, Thibault Damour, Holmfridur Hannesdottir, Johannes Henn, Enrico Herrmann, Gary Horowitz, Gregory Korchemsky, Yao Ma, Sebastian Mizera, Michael Ruf, Lorenzo Tancredi, Gautam Satishchandran, George Sterman, Robert Szafron, and Céline Zwikel for insightful discussions. AH is grateful to the Simons Foundation as well as the Edward and Kiyomi Baird Founders’ Circle Member Recognition for their support. SDA was supported by the European Research Council, under grant ERC-AdG-885414. R.R. is supported by the U.S. Department of Energy (DOE) under award number DE-SC00019066 and by a Senior Fellowship at the Institute for Theoretical Studies, ETH Zürich. F.T. is supported by the startup grant from the department of physics at Fudan University. Part of the work reported in this paper was carried out while R.R. and F.T. participated in the “What is Particle Theory?” program at the Kavli Institute for Theoretical Physics (KITP) and supported in part by grant NSF PHY-2309135.

### v2

@@ 5 Conclusions
We contrast here two complementary observations. On the one hand, the transformation of the NP scalars under general coordinates transformations, large or small, suggests that their leading order in the $1/|\vec{x}|$ expansion cannot be changed by diffeomorphisms that preserves the asymptotic-Minkowski nature of the spacetime. In particular, the leading-order peeling violation cannot be removed by such transformations. On the other, a BMS transformation with parameter twice that of Veneziano and Vilkovisky Veneziano:2022zwh flips the sign of the disconnected-cut contribution to the waveform Bini:2024rsy ; Elkhidir:2024izo and, in an off-shell regularization Elkhidir:2024izo , appears to completely cancel the tail-related logarithmic terms in Eq. (2.2), $iG(p_{1}\cdot k\ln(u_{1}\cdot k)^{2}/k^{2})+p_{2}\cdot k\ln(u_{2}\cdot k)^{2}/k^{2}))$. In this BMS frame, which correspond to a very specific choice of boundary conditions at finite time and is distinct from both the canonical and the intrinsic ones, the novel peeling violation that we identified in Sec 4.4 appears to cancel out. From our analysis, it is unclear why such choice of boundary conditions are special and we leave further investigations for future work. In particular, it would be interesting to explore the precise map relating different BMS frames at $\mathcal{I}^{-}$ to boundary conditions at space-like hypersurfaces at $t\rightarrow$-$\infty$.
It would be fascinating if the subleading components of the metric in $|\vec{x}|^{-1}$ were experimentally measurable. Naively, these effects are enormously suppressed by the vast distance between astrophysical events and gravitational-wave observatories, making them largely formal considerations. Nevertheless, the computational techniques developed in this paper could prove useful in contexts where finite-distance effects become relevant, such as in comparisons with numerical relativity simulations Mitman:2024uss and at colliders Collins:2019ozc .
The framework developed in this paper can be used to compute the scattering metric throughout spacetime. This calculation amounts to evaluating exactly the Fourier transform over the external-graviton momentum of the final-state graviton one-point function. Assuming that such an exact metric were available, it should be possible to direct test of asymptotic simplicity by computing the conformal structure and demonstrating that they are not regular at null infinity.
@@ Acknowledgements.
We thank Martin Beneke, Donato Bini, Thibault Damour, Holmfridur Hannesdottir, Carlo Heissenberg, Johannes Henn, Enrico Herrmann, Gary Horowitz, Gregory Korchemsky, Yao Ma, Sebastian Mizera, Michael Ruf, Lorenzo Tancredi, Gautam Satishchandran, George Sterman, Robert Szafron, and Céline Zwikel for insightful discussions. AH is grateful to the Simons Foundation as well as the Edward and Kiyomi Baird Founders’ Circle Member Recognition for their support. SDA was supported by the European Research Council, under grant ERC-AdG-885414. R.R. is supported by the U.S. Department of Energy (DOE) under award number DE-SC00019066 and by a Senior Fellowship at the Institute for Theoretical Studies, ETH Zürich. F.T. is supported by the startup grant from the department of physics at Fudan University. Part of the work reported in this paper was carried out while R.R. and F.T. participated in the “What is Particle Theory?” program at the Kavli Institute for Theoretical Physics (KITP) and supported in part by grant NSF PHY-2309135.